# Student reasoning about quantum mechanics while working with physical experiments

### v1

@@ VI Conclusions
We first identified four resources students activated while engaging with the idea that single photons can interfere with themselves. Both pairs of students activated the resource Knowledge of classical wave interference, one pair also activated the resource Knowledge of wave functions, and the other pair instead activated the resources Probabilities represent possible quantum outcomes and Thinking in terms of information. These resources and the ways students used them provide an example for instructors of the background knowledge students might need as they work with the single-photon experiments, as well as ways instructors may be able to help students build off of their already existing ideas.

### v2

@@ VI Conclusions
We first identified several resources students activated while engaging with the idea that single photons can interfere with themselves. Both pairs of students activated various resources related to classical wave interference, one pair also activated several resources related to wave functions, and the other pair instead activated the resources Quantum outcomes are probabilistic and Thinking in terms of information. These resources and the ways students used them provide an example for instructors of the background knowledge students might need as they work with the single-photon experiments, as well as ways instructors may be able to help students build off of their already existing ideas.